OpenStudy (anonymous):

Write an equation that is perpendicular to y=(3/2)x+2 and crosses the point (2,5)? I tried this one myself but got the wrong answer. :/

OpenStudy (anonymous):

With the point slop formula i got.. y=(3/2)x+2 and that can't be right.

OpenStudy (anonymous):

*slope haha

OpenStudy (anonymous):

The product of the slopes of two perpendicular lines must be -1 So, in the ecuation y=(3/2)x+2, the slope us 3/2. So, let m be the slope of the requested perpendicular line, then (3/2)*m = -1 Then you solve for m, and plugin in to the point slope formula.

OpenStudy (anonymous):

so the slope for the other one has to be the opposite inverse, i think. it'd be 2/3?

OpenStudy (anonymous):

negative 2/3, oops

OpenStudy (anonymous):

Correct the slope is -2/3, then you can plug the numbers in the slope point ecuation y-y_1 = m(x-x1) Where (x1,y1) is the point adn m the slope.

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

Welcome